# Model-guided Fuzzing of Distributed Systems

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software testing and debuggingccs: Theory of computation Distributed computing models
@@ 24 Buffer buffer;
    Register(w), Register(t), Request(r), Execute(r),
    Terminate(w), Flush
@@
    Register(t), Register(w), Request(r), Terminate(w),
    Execute(r), Flush
@@
    Register(t), Register(w), Request(r),
    Terminate(w), Flush, Execute(r)
@@ 3.2. Coverage-guided Fuzzing of Distributed Systems
4
5proc $\mathtt{coverageGuidedFuzzer}(S,M,T_{0})$
@@ 15                  $totalCoverage\leftarrow totalCoverage\cup coverage$
18
19proc $\mathtt{executeAndGetCoverage}(S,M,t)$
@@ 4. Implementation of ModelFuzz
  • The protocol actions for processing protocol messages, ClientRequest, HandleRequestVoteRequest, HandleRequest- VoteResponse, HandleAppendEntriesRequest, HandleAppend- EntriesRequest, HandleNilAppendEntriesResponse, identified with protocol arguments (e.g., term and index numbers): The system events of exchanging these messages with corresponding arguments map to the actions.
@@ 5.1. Microbenchmark in Coyote
Figure 3. Testing the microbenchmark. Test coverage for $m=6$ workers and $n=40$ tasks
Figure 4. Testing the microbenchmark. #tests to find the bug for varying #tasks with $m=6$ workers.
@@
Figure 3 shows the coverage of the abstract states of the microbenchmark with $m=6$ workers and $n=40$ tasks, which is representative of different parameter configurations. Since the microbenchmark is a small example with a small state space, the difference in the explored number of unique abstract states among different testing approaches is not large. However, the results show ModelFuzz’s ability to cover more abstract states compared to random testing and trace coverage guidance. Our Mann-Whitney U-tests show that ModelFuzz achieves statistically significantly better coverage results at $\alpha=0.05$, compared to random testing and trace-guided fuzzing with p-values ${\{0.0001,0.0004\}}$.
@@
We observe a trend with ModelFuzz where it consistently detects the bug faster than random and trace coverage guided testing approaches with increasing $m$ and $n$. Figure 4 plots the number of test iterations to find the bug for increasing $n$ number of task messages with a fixed $m=6$ processes. The results show that the increasing number of task messages makes the bug more difficult to detect as it is triggered deep in the execution space. This effect is most evidently seen with pure random testing, as it fails to detect the bug after $n=10$. Trace coverage guided testing achieves a more consistent variance among different campaigns. However, we observe a trend in its median value for the first iteration to detect the bug, where it declines as the bug gets harder to detect with increasing $n$. The performance degradation is not as significant for ModelFuzz, as its median value does not change drastically among the experiments.
@@ 5.2. Etcd-raft
Figure 5. Testing etcd. Test coverage of abstract states
@@
Table 2. Testing etcd. Average number of test iterations to find the synthetic bug. The new bug is only detected by ModelFuzz.
@@
Figure 5 reports the test coverage of the test harnesses in the number of abstract states observed with different strategies. The results show that model-guided test generation of ModelFuzz outperforms unguided random testing (Random), coverage-guided fuzzing using line coverage (Line), and coverage-guided fuzzing using trace coverage (Trace) in the explored number of unique system states. ModelFuzz covers 1.22x more states than Random, 1.23x more states than Line, and 1.48x more states than Trace coverage approaches. Comparing model-guided fuzzing and structural guidance, we find that in both cases, the code coverage saturates at $47.9\%$. Similarly, comparing model-guided fuzzing and trace-based fuzzing, we find that in both cases, we explore 10k unique traces. To mitigate the randomness, we perform Mann-Whitney U Test and conclude that ModelFuzz’s coverage of model states is significantly higher than all other approaches, with p values of $1.53\mathrm{e}{-10}$.
@@
To answer RQ2, we analyze the number of detected bugs and the number of test iterations to discover a bug using different strategies. While the seeded bug can be found in each of the 20 trials by all of the strategies, the new bug can only be found using ModelFuzz. Figure 2 reports the average number of iterations to discover the seeded bug using different strategies. The bug can be found by random search faster on average. This can be explained by the characteristics of the bug, which is easily triggered in executions with quorums of only $n/3+1$ processes.
@@ 5.3. RedisRaft
Figure 6 illustrates the average coverage measures for 20K test iterations for each testing strategy. Similar to Etcd, we show that ModelFuzz can obtain better coverage over model states compared to random testing, line coverage guided, and trace coverage-guided fuzzing strategies. On average, ModelFuzz observes 2.58x more states than random exploration, 2.43x more than line coverage, and 2.84x more coverage than trace-guided fuzzing. As with Etcd-Raft, we answer  RQ1 with an observation that model guidance outperforms random exploration and other coverage-guidance strategies in the coverage of explored system states.
@@
We also analyze the branch coverage of the tests and observe that guiding the test executions using model coverage does not degrade the coverage over traditional code coverage metrics. Table 4 reports the mean and standard deviation branch coverage of the different guidance methods. We measure branch coverage of the source code in C using the gcov tool. However, we observe high variance in the data, which we attribute to the coverage measurement tool. As we run multiple copies of the source code during each test iteration, the branch coverage is combined for each copy. However, we found that the gcov tool does not always merge the coverage values of concurrent invocations correctly. Failing to merge coverage information results in the high variance we observe with the branch coverage measures presented in the table.
Figure 6. Testing RedisRaft. Test coverage of abstract states
@@
Table 4. Testing RedisRaft. Average branch coverage of 20 runs.
@@ 8 Process crashes when updating the snapshot index offset
Table 6. The number of RedisRaft tests for the first occurrences of the bugs using different guidance strategies.
@@
Table 7. Pairwise $\hat{A}_{12}$ statistic results against ModelFuzz.
Among all 14 bugs, ModelFuzz found more bugs than other guidance approaches. Specifically, ModelFuzz found 13 bugs, while random exploration, trace-guided, and line-guided found only 10 of them. Furthermore, ModelFuzz is faster in reproducing 7 of the bugs compared to other approaches, where Random exploration and Line-guided each find three, and trace-guided finds one bug faster. For each bug and guidance method, the table in Figure 6 lists the number of runs that find the bug (in parenthesis) and the average first occurrence of the bug in the successful runs. For each bug, we highlight the lowest average first occurrence in bold. Among the bugs, three of them are found only using ModelFuzz. We classify a bug as rare if it was found in at most five trials. ModelFuzz finds four of the rare bugs faster than the other approaches.
For each of the 14 bugs, we calculate Vargha and Delaney’s $\hat{A}_{12}$ statistics to analyze the pair-wise statistical significance of the results. Table 7 reports the $\hat{A}_{12}$ statistics against ModelFuzz for each bug, highlighting the results with statistical significance in bold. The analysis shows that ModelFuzz detects the bugs ${\{6,12,13,14\}}$ statistically significantly faster than all other guidance approaches, ${\{2,7,9,10\}}$ statistically significantly faster than some of the approaches, and comparably faster for the remainder. While this indicates its ability to detect bugs faster, the $\hat{A}_{12}$ results do not draw clear conclusions on the statistical significance.
@@
Figure 7 reports the mean, standard deviation, and confidence intervals for all approaches, and Figure 8 illustrates the Tukey HSD results. We can observe that the mean iteration to detect a bug for ModelFuzz lies outside of the other approaches. This suggests that the differences between ModelFuzz and all of the remaining approaches are statistically significant, whereas there is no significant difference between random exploration, line coverage, and trace coverage guidance. Note that we perform ANOVA and post-hoc Tukey HSD tests only for RedisRaft, as we analyze numerous bugs with which we could form a sample set.
@@
Figure 7. Mean, standard deviation, lower confidence ($CI_{l}$), and upper confidence interval values ($CI_{u}$) of the first iteration to find a bug.
Figure 8. The Tukey HSD test plot.

### v2

@@ Abstract.
copyright: acmlicenseddoi: XXXXXXX.XXXXXXXccs: Software and its engineering Software testing and debuggingccs: Theory of computation Distributed computing models
@@ 24 Buffer buffer;
    Register(w), Register(t), Request(r), Execute(r), Terminate(w), Flush
@@
    Register(t), Register(w), Request(r), Terminate(w), Execute(r), Flush
@@
    Register(t), Register(w), Request(r), Terminate(w), Flush, Execute(r)
@@ 3.2. Coverage-guided Fuzzing of Distributed Systems
3
4proc $\mathtt{coverageGuidedFuzzer}(S,M,T_{0})$
@@ 14                  $totalCoverage\leftarrow totalCoverage\cup coverage$
17
18proc $\mathtt{executeAndGetCoverage}(S,M,t)$
@@ 4. Implementation of ModelFuzz
  • The protocol actions for processing protocol messages,
    ClientRequest, HandleRequestVoteRequest, HandleRequest- VoteResponse,
    HandleAppendEntriesRequest, HandleAppend- EntriesRequest, HandleNilAppendEntriesResponse,
    identified with protocol arguments (e.g., term and index numbers): The system events of exchanging these messages with corresponding arguments map to the actions.
@@ 5.1. Microbenchmark in Coyote
Figure 3. Testing the microbenchmark. (a) Test coverage for $m=6$ workers and $n=40$ tasks (b) #tests to find the bug for varying #tasks with $m=6$ workers.
@@
Figure 3(a) shows the coverage of the abstract states of the microbenchmark with $m=6$ workers and $n=40$ tasks, which is representative of different parameter configurations. Since the microbenchmark is a small example with a small state space, the difference in the explored number of unique abstract states among different testing approaches is not large. However, the results show ModelFuzz’s ability to cover more abstract states compared to random testing and trace coverage guidance. Our Mann-Whitney U-tests show that ModelFuzz achieves statistically significantly better coverage results at $\alpha=0.05$, compared to random testing and trace-guided fuzzing with p-values ${\{0.0001,0.0004\}}$.
@@
We observe a trend with ModelFuzz where it consistently detects the bug faster than random and trace coverage guided testing approaches with increasing $m$ and $n$. Figure 3(b) plots the number of test iterations to find the bug for increasing $n$ number of task messages with a fixed $m=6$ processes. The results show that the increasing number of task messages makes the bug more difficult to detect as it is triggered deep in the execution space. This effect is most evidently seen with pure random testing, as it fails to detect the bug after $n=10$. Trace coverage guided testing achieves a more consistent variance among different campaigns. However, we observe a trend in its median value for the first iteration to detect the bug, where it declines as the bug gets harder to detect with increasing $n$. The performance degradation is not as significant for ModelFuzz, as its median value does not change drastically among the experiments.
@@ 5.2. Etcd-raft
Figure 4. Testing etcd. (a) Test coverage of abstract states (b) Average number of test iterations to find the synthetic bug. The new bug is only detected by ModelFuzz.
@@
Figure LABEL:fig:etcd-cov-1 reports the test coverage of the test harnesses in the number of abstract states observed with different strategies. The results show that model-guided test generation of ModelFuzz outperforms unguided random testing (Random), coverage-guided fuzzing using line coverage (Line), and coverage-guided fuzzing using trace coverage (Trace) in the explored number of unique system states. ModelFuzz covers 1.22x more states than Random, 1.23x more states than Line, and 1.48x more states than Trace coverage approaches. Comparing model-guided fuzzing and structural guidance, we find that in both cases, the code coverage saturates at $47.9\%$. Similarly, comparing model-guided fuzzing and trace-based fuzzing, we find that in both cases, we explore 10k unique traces. To mitigate the randomness, we perform Mann-Whitney U Test and conclude that ModelFuzz’s coverage of model states is significantly higher than all other approaches, with p values of $1.53\mathrm{e}{-10}$.
@@
To answer RQ2, we analyze the number of detected bugs and the number of test iterations to discover a bug using different strategies. While the seeded bug can be found in each of the 20 trials by all of the strategies, the new bug can only be found using ModelFuzz. Figure LABEL:tbl:etcd-bugs reports the average number of iterations to discover the seeded bug using different strategies. The bug can be found by random search faster on average. This can be explained by the characteristics of the bug, which is easily triggered in executions with quorums of only $n/3+1$ processes.
@@ 5.3. RedisRaft
Figure LABEL:fig:redisraft-cov illustrates the average coverage measures for 20K test iterations for each testing strategy. Similar to Etcd, we show that ModelFuzz can obtain better coverage over model states compared to random testing, line coverage guided, and trace coverage-guided fuzzing strategies. On average, ModelFuzz observes 2.58x more states than random exploration, 2.43x more than line coverage, and 2.84x more coverage than trace-guided fuzzing. As with Etcd-Raft, we answer  RQ1 with an observation that model guidance outperforms random exploration and other coverage-guidance strategies in the coverage of explored system states.
@@
We also analyze the branch coverage of the tests and observe that guiding the test executions using model coverage does not degrade the coverage over traditional code coverage metrics. Table LABEL:tbl:redis-branch-cov reports the mean and standard deviation branch coverage of the different guidance methods. We measure branch coverage of the source code in C using the gcov tool. However, we observe high variance in the data, which we attribute to the coverage measurement tool. As we run multiple copies of the source code during each test iteration, the branch coverage is combined for each copy. However, we found that the gcov tool does not always merge the coverage values of concurrent invocations correctly. Failing to merge coverage information results in the high variance we observe with the branch coverage measures presented in the table.
@@
Figure 5. Testing RedisRaft. (a) Test coverage of abstract states (b) Average branch coverage of 20 runs.
@@ 8 Process crashes when updating the snapshot index offset
Figure 6. (a) The number of RedisRaft tests for the first occurrences of the bugs using different guidance strategies. (b) Pairwise $\hat{A}_{12}$ statistic results against ModelFuzz.
Among all 14 bugs, ModelFuzz found more bugs than other guidance approaches. Specifically, ModelFuzz found 13 bugs, while random exploration, trace-guided, and line-guided found only 10 of them. Furthermore, ModelFuzz is faster in reproducing 7 of the bugs compared to other approaches, where Random exploration and Line-guided each find three, and trace-guided finds one bug faster. For each bug and guidance method, the table in Figure LABEL:tbl:redis-bugs-alt lists the number of runs that find the bug (in parenthesis) and the average first occurrence of the bug in the successful runs. For each bug, we highlight the lowest average first occurrence in bold. Among the bugs, three of them are found only using ModelFuzz. We classify a bug as rare if it was found in at most five trials. ModelFuzz finds four of the rare bugs faster than the other approaches.
For each of the 14 bugs, we calculate Vargha and Delaney’s $\hat{A}_{12}$ statistics to analyze the pair-wise statistical significance of the results. Table LABEL:tbl:redis-bug-stats reports the $\hat{A}_{12}$ statistics against ModelFuzz for each bug, highlighting the results with statistical significance in bold. The analysis shows that ModelFuzz detects the bugs ${\{6,12,13,14\}}$ statistically significantly faster than all other guidance approaches, ${\{2,7,9,10\}}$ statistically significantly faster than some of the approaches, and comparably faster for the remainder. While this indicates its ability to detect bugs faster, the $\hat{A}_{12}$ results do not draw clear conclusions on the statistical significance.
@@
Figure 7(a) reports the mean, standard deviation, and confidence intervals for all approaches, and Figure 7(b) illustrates the Tukey HSD results. We can observe that the mean iteration to detect a bug for ModelFuzz lies outside of the other approaches. This suggests that the differences between ModelFuzz and all of the remaining approaches are statistically significant, whereas there is no significant difference between random exploration, line coverage, and trace coverage guidance. Note that we perform ANOVA and post-hoc Tukey HSD tests only for RedisRaft, as we analyze numerous bugs with which we could form a sample set.
@@
Figure 7. (a) Mean, standard deviation, lower confidence ($CI_{l}$), and upper confidence interval values ($CI_{u}$) of the first iteration to find a bug. (b) The Tukey HSD test plot.